Silver Company needs $1,000,000 in order to buy a new equipment for its business. Silver’s net income is $950,000. Their target capital structure is D/E=1/3. How much Silver can pay out as dividends?  

Select one:

a. $220,000

b. $200,000

c. $210,000

d. $230,000
Answer #1
  • The company needs to maintain the target capital structure represented by Debt/Equity ratio of 1/3 even after investing in new equipment.
  • D/E = Debt / Equity = 1/3
  • Amount of equity funding for the new equipment = $1,000,000 × 3/(1+3) = $750,000
  • Therefore, $750,000 of equity is required to maintain the target D/E ratio which is sourced from the retained earnings or net income.
  • Silver can pay out as dividends: Net income - Amount required for investment = $950,000 - $750,000 = $200,000
  • $250,000 will be financed by debt.
  • Now D/E ratio of the new investment = Debt/Equity = $250,000/$750,000 = 1/3